# Kernel-based optimization of measurement operators for quantum reservoir computers

### v1

@@ Appendix B Generalization to multiple outputs/classes
For multi-class classification with labels $y_{i}\in\{1,\ldots,C\}$, one can use one-hot targets and interpret the components $f_{c}^{*}(\mathbf{x})$ as class scores, predicting $\hat{y}(\mathbf{x})=\arg\max_{c}f_{c}^{*}(\mathbf{x})$ (or applying a softmax to obtain class probabilities). One thus replaces the single observable $M$ by $C$ observables $\{M_{c}\}_{c=1}^{C}$ (equivalently a coefficient matrix $\mathbf{A}\in\mathbb{R}^{P\times C}$) with regularizer $\frac{\lambda}{2}\sum_{c}(|M_{c}|^{\text{HS}})^{2}$. The representer form then holds component-wise as $M_{c}^{*}=\sum_{i}\alpha_{ic}^{*}\rho(\mathbf{x}_{i})$, and one-vs-rest hinge/logistic training reduces to $C$ standard kernel problems with the same Gram matrix $\mathbf{K}$.
@@ Appendix C Diagonalization
$\beta_{\mathbf{s}}=\mathrm{tr}(\Lambda Z_{\mathbf{s}})=2^{-N}\sum_{\mathbf{b}\in\{0,1\}^{N}}\lambda_{\mathbf{b}}\,(-1)^{\mathbf{s}\cdot\mathbf{b}}.$ (54)
@@ Appendix D Pauli basis truncation
Assume we measure only a subset of Pauli observables in the expansion of $M^{*}$ [Eq. 13]. Let the corresponding index set be $S\subset{1,\ldots,d^{2}}$ and define the truncated observable $M_{\text{trunc}}=\sum_{k\in S}c_{k}^{*}\mathcal{O}_{k}$. By Eq. 2, the pointwise prediction error on any input $\mathbf{x}$ is $\Delta f(\mathbf{x})=\mathrm{tr}\big((M^{*}-M_{\text{trunc}})\rho(\mathbf{x})\big)$. Using the Cauchy–Schwarz inequality in the Hilbert–Schmidt inner product yields $|\Delta f(\mathbf{x})|\leq|M^{*}-M_{\text{trunc}}|^{\text{HS}}\,|\rho(\mathbf{x})|^{\text{HS}}$. Since for pure feature states $|\rho(\mathbf{x})|^{\text{HS}}=1$, the worst-case prediction error is set by the magnitude of the omitted coefficients:
$|\Delta f(\mathbf{x})|\leq(|M^{*}-M_{\text{trunc}}|^{\text{HS}})^{2}=\sum_{k\notin S}(c_{k}^{*})^{2},$ (56)
where we used orthonormality of the basis.
@@ Appendix E Classical ELM
$f(\mathbf{x})=\phi(\mathbf{x})^{T}\mathbf{w}$ (57)
Recall that a QELM can be mapped to this form by identifying $\phi_{k}(\mathbf{x})=\mathrm{tr}[M_{k}\rho(\mathbf{x})]$ for a set of $m$ observables $\{M_{k}\}$ [see Eq. 51]. Assuming a least-squares loss and $P$ samples
$\mathcal{L}(\{\mathbf{x}_{j},y_{j}\})=\frac{1}{2}\sum_{j=1}^{P}\left[y_{j}-f(\mathbf{x}_{j})\right]^{2}+\frac{\lambda}{2}\|\mathbf{w}\|^{2}$ (58)
the optimal solution follows as
@@
$k(\mathbf{x},\mathbf{x}^{\prime})=\phi(\mathbf{x})^{T}\phi(\mathbf{x}^{\prime})\in\mathbb{R}$ (60)
@@
$f(\mathbf{x})=\phi(\mathbf{x})^{T}\mathbf{w}^{*}=\phi(\mathbf{x})^{T}\Phi^{T}\boldsymbol{\alpha}^{*}=\sum_{j=1}^{P}\phi(\mathbf{x})^{T}\phi(\mathbf{x}_{j})\alpha_{j}=\sum_{j=1}^{P}\alpha_{j}k(\mathbf{x},\mathbf{x}_{j})$ (63)
@@ Appendix F QRC with stateful reservoir
We describe here in some more detail how the kernel-based optimization can be applied to the case of a QRC with internal memory, complementing Appendix A and demonstrating the viability of prescription Eq. 3. Let the input register of the QRC have $N_{I}$ qubits and the memory register $N_{A}$ qubits. At time $t$, a classical input $\mathbf{z}_{t}\in\mathbb{R}^{d^{\prime}}$ (typically $d^{\prime}=N_{I}$) is encoded via the density matrix $\eta_{I}(\mathbf{z}_{t})$. The joint state $\eta_{t}=\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)$ is then time-evolved by a unitary $U$, giving
$\rho(t)\equiv\eta_{t+1}=U\eta_{t}U^{\dagger},$ (64)
which corresponds to the QELM feature map in Eq. 2. Subsequently, the input is measured/reset and discarded, and the memory alone is carried to the next step, described by the CPTP map
$\eta_{M}(t+1)=\mathrm{tr}_{I}\!\left[\,U\bigl(\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)\bigr)U^{\dagger}\right].$
@@
$(\mathbf{z}_{0},\dots,\mathbf{z}_{t})=:x_{t}\longmapsto\rho(x_{t})\in\mathcal{M}(\mathcal{H}).$ (65)
Due to the fading memory of the reservoir, the dependence on past inputs $\mathbf{z}_{0},\dots,\mathbf{z}_{t-1}$ is effectively cut off at a certain finite depth $L$. Given an reservoir equilibration time $t_{\mathrm{eq}}$, we thus define the training index set
$\mathcal{T}_{\mathrm{train}}:=\{t_{\mathrm{eq}},\dots,T-1\},\qquad P:=|\mathcal{T}_{\mathrm{train}}|,$ (66)
and an effective input history (or window)
$x_{t}:=(\mathbf{z}_{t-L},\ldots,\mathbf{z}_{t})\in(\mathbb{R}^{d^{\prime}})^{L+1},$ (67)
for $t\in\mathcal{T}_{\mathrm{train}}$. Let $\rho(x_{t})$ denote the reservoir state obtained from driving the reservoir with that history from some fixed initialization (the precise initialization is irrelevant by assumption). The corresponding readout is
@@
For targets $\{y_{t}\}_{t=0}^{T}$, we consider the optimization problem
$\min_{M\in\mathcal{M}(\mathcal{H})}\frac{1}{2}\sum_{t\in\mathcal{T}_{\mathrm{train}}}\bigl(\mathrm{tr}(M\rho(x_{t}))-y_{t}\bigr)^{2}+\frac{\lambda}{2}\mathrm{tr}(M^{2}),$ (69)
@@
with $\boldsymbol{\alpha}^{*}=(\alpha_{t}^{*})_{t\in\mathcal{T}_{\mathrm{train}}}$ and the target vector $\mathbf{y}=(y_{t})_{t\in\mathcal{T}_{\mathrm{train}}}\in\mathbb{R}^{P}$. For any test history/window $x$ (obtained from any test sequence driven through the same reservoir),
$f^{*}(x)=\sum_{t\in\mathcal{T}_{\mathrm{train}}}\alpha_{t}^{*}K(x_{t},x).$ (73)
For a fixed driving sequence, one may write $K(t,t^{\prime})$ as shorthand for $K(x_{t},x_{t^{\prime}})$. This is the direct stateful counterpart of the kernel representation in Eq. 11, with the feature map $\rho(x)$ replaced by the stateful reservoir map $(z_{t-L},\dots,z_{t})=:\mathbf{x}_{t}\mapsto\rho(t)$. For vector-valued targets $y_{t}\in\mathbb{R}^{C}$ (for example, when predicting all components of $z_{t+1}$), one can apply the same construction component-wise, as in the stateless case discussed below Eq. 10.
@@ Appendix G Product state encodings and measurements
The following table summarizes the single-qubit expectation values obtained for rotational and amplitude encoding:
@@
This implies, in particular, that measuring any product state with a Pauli string $P$ containing a $Y$-factor will always yield zero. Action of a unitary breaks Eq. 74 as it typically spreads any $\sigma^{p}$ into a mixture of all other Paulis, such that measuring a $Y$ then will result in nonzero contributions.

### v2

@@ Appendix B Generalization to multiple outputs/classes
For multi-class classification with labels $y_{i}\in\{1,\ldots,C\}$, one can use one-hot targets and interpret the components $f_{c}^{*}(\mathbf{x})$ as class scores, predicting $\hat{y}(\mathbf{x})=\arg\max_{c}f_{c}^{*}(\mathbf{x})$ (or applying a softmax to obtain class scores). One thus replaces the single observable $M$ by $C$ observables $\{M_{c}\}_{c=1}^{C}$ (equivalently a coefficient matrix $\mathbf{A}\in\mathbb{R}^{P\times C}$) with regularizer $\frac{\lambda}{2}\sum_{c}(|M_{c}|^{\text{HS}})^{2}$. The representer form then holds component-wise as $M_{c}^{*}=\sum_{i}\alpha_{ic}^{*}\rho(\mathbf{x}_{i})$, and one-vs-rest hinge/logistic training reduces to $C$ standard kernel problems with the same Gram matrix $\mathbf{K}$.
@@ Appendix C Diagonalization
$\beta_{\mathbf{s}}=2^{-N}\mathrm{tr}(\Lambda Z_{\mathbf{s}})=2^{-N}\sum_{\mathbf{b}\in\{0,1\}^{N}}\lambda_{\mathbf{b}}\,(-1)^{\mathbf{s}\cdot\mathbf{b}}.$ (54)
@@ Appendix D Pauli basis truncation
Assume we measure only a subset of Pauli observables in the expansion of $M^{*}$ [Eq. 13]. Let the corresponding index set be $S\subset\{0,\ldots,D^{2}-1\}$ and define the truncated observable $M_{\text{trunc}}=\sum_{k\in S}c_{k}P_{k}$. By Eq. 39, the pointwise prediction error on any input $\mathbf{x}$ is $\Delta f(\mathbf{x})=\mathrm{tr}\big((M^{*}-M_{\text{trunc}})\rho(\mathbf{x})\big)$. Using the Cauchy–Schwarz inequality in the Hilbert–Schmidt inner product yields $|\Delta f(\mathbf{x})|\leq|M^{*}-M_{\text{trunc}}|^{\text{HS}}\,|\rho(\mathbf{x})|^{\text{HS}}$. Since for any feature state $|\rho(\mathbf{x})|^{\text{HS}}\leq 1$, the worst-case prediction error is bounded by the magnitude of the omitted coefficients:
$|\Delta f(\mathbf{x})|\leq\|M^{*}-M_{\text{trunc}}\|_{\text{HS}}=\Big(2^{N}\sum_{k\notin S}c_{k}^{2}\Big)^{1/2},$ (56)
where we used orthogonality of the basis.
@@ Appendix E Classical ELM
$f(\mathbf{x})=\phi(\mathbf{x})^{T}\mathbf{w}.$ (57)
Recall that a QELM can be mapped to this form by identifying $\phi_{k}(\mathbf{x})=\mathrm{tr}[M_{k}\rho(\mathbf{x})]$ for a set of $m$ observables $\{M_{k}\}$ [see Eq. 51]. Assuming a least-squares loss,
$\mathcal{L}(\{\mathbf{x}_{j},y_{j}\})=\frac{1}{2}\sum_{j=1}^{P}\left[y_{j}-f(\mathbf{x}_{j})\right]^{2}+\frac{\lambda}{2}\|\mathbf{w}\|^{2},$ (58)
and $P$ samples, the optimal solution follows as
@@
$k(\mathbf{x},\mathbf{x}^{\prime})=\phi(\mathbf{x})^{T}\phi(\mathbf{x}^{\prime})\in\mathbb{R},$ (60)
@@
$f(\mathbf{x})=\phi(\mathbf{x})^{T}\mathbf{w}^{*}=\phi(\mathbf{x})^{T}\Phi^{T}\boldsymbol{\alpha}^{*}=\sum_{j=1}^{P}\phi(\mathbf{x})^{T}\phi(\mathbf{x}_{j})\alpha_{j}=\sum_{j=1}^{P}\alpha_{j}k(\mathbf{x},\mathbf{x}_{j}).$ (63)
@@ Appendix F QRC with stateful reservoir
We describe here in some more detail how the kernel-based optimization can be applied to the case of a QRC with internal memory, complementing Appendix A and demonstrating the viability of prescription Eq. 3. Let the input register of the QRC have $N_{I}$ qubits and the memory register $N_{A}$ qubits. At time $t$, a classical input $\mathbf{z}_{t}\in\mathbb{R}^{d^{\prime}}$ (typically $d^{\prime}=N_{I}$) is encoded via the density matrix $\eta_{I}(\mathbf{z}_{t})$. The joint state $\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)$ is time-evolved by a unitary $U$, giving
$\eta(t)=U\left(\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)\right)U^{\dagger}.$ (64)
Subsequently, the input is measured/reset and discarded, and the memory alone is carried to the next step, described by the CPTP map
$\eta_{M}(t+1)=\mathrm{tr}_{I}\!\left[\eta(t)\right].$ (65)
@@
$(\mathbf{z}_{0},\dots,\mathbf{z}_{t})=:x_{t}\longmapsto\rho(x_{t})=\eta(t)\in\mathcal{M}(\mathcal{H}),$ (66)
where $\rho(x_{t})$ is obtained after sequentially processing the $\{\mathbf{z}_{t^{\prime}}\}$ via Eqs. 64 and 65. Due to the fading memory of the reservoir, the dependence on inputs earlier than time tLt-L is negligible. We assume that any transient associated with the reservoir initialization has already decayed before the observed sequence begins, e.g. by driving the reservoir for times $t<0$. We therefore work, instead of (66), with the effective input history (window)
$x_{t}:=(\mathbf{z}_{t-L},\ldots,\mathbf{z}_{t})\in(\mathbb{R}^{d^{\prime}})^{L+1},\qquad t=L,\ldots,T-1,$ (67)
and define P=TLP=T-L training samples. Let $\rho(x_{t})$ denote the reservoir state obtained from driving the reservoir with that history from some fixed initialization (the precise initialization is irrelevant by assumption). The corresponding readout is
@@
For targets $\{y_{t}\}_{t=L}^{T-1}$, we consider the optimization problem
$\min_{M\in\mathcal{M}(\mathcal{H})}\frac{1}{2}\sum_{t=L}^{T-1}\bigl(\mathrm{tr}(M\rho(x_{t}))-y_{t}\bigr)^{2}+\frac{\lambda}{2}\mathrm{tr}(M^{2}),$ (69)
@@
with $\boldsymbol{\alpha}^{*}=(\alpha_{t}^{*})_{t=L}^{T-1}$ and the target vector $\mathbf{y}=(y_{t})_{t=L}^{T-1}\in\mathbb{R}^{P}$. For any test history/window $x$ (obtained from any test sequence driven through the same reservoir),
$f^{*}(x)=\sum_{t=L}^{T-1}\alpha_{t}^{*}K(x_{t},x).$ (73)
For a fixed driving sequence, one may write $K(t,t^{\prime})$ as shorthand for $K(x_{t},x_{t^{\prime}})$. This is the direct stateful counterpart of the kernel representation in Eq. 11, with the feature map $\rho(x)$ replaced by the stateful reservoir map $(\mathbf{z}_{t-L},\dots,\mathbf{z}_{t})=:x_{t}\mapsto\rho(x_{t})$. For vector-valued targets $\mathbf{y}_{t}\in\mathbb{R}^{C}$ (for example, when predicting all components of $\mathbf{z}_{t+1}$), one can apply the same construction component-wise, as in the stateless case discussed below Eq. 10.
@@ Appendix G Product state encodings and measurements
Table 4 summarizes the single-qubit expectation values obtained for rotational and amplitude encoding.
@@
Table 4: Single-qubit expectation values for rotational and amplitude encoding.
This implies, in particular, that measuring the above encoded product states with a Pauli string $P$ containing a $Y$-factor will always yield zero  [^5] This does not apply to the dense encoding in Eq. 27.. Action of a unitary breaks Eq. 74 as it typically spreads any $\sigma^{p}$ into a mixture of all other Paulis, such that measuring a $Y$ then will result in nonzero contributions.